# The Physics of Preference: Unravelling Imprecision of Human Preferences through Magnetisation Dynamics

### v1

@@ Abstract
Using the numerical method described below, in Fig. 1b we plot the dynamics of the $M_{z}$ component of the magnetisation vector ${\bf M}$ in the “free” layer as a function of the electric current strength for different value of the applied magnetic field $H_{0}$. The trajectories of ${\bf M}$ for $H_{0}=8$ kOe computed at the simulated time of 1 ns are shown in Fig. 3b–c. Figure 1b demonstrates that the magnetisation dynamics correctly approximates the sigmoid probabilistic PR curve in Fig. 1a but the strength of the driving electric current plays the role of parameter $A$. The experimental curves in Fig. 2 can be regarded as the particular cases of the sigmoid curve and they can be approximated using magnetisation dynamics at different values of $H_{0}$, which is demonstrated below.
@@
Study 2—We validate our model using a dataset obtained from a “Deal or No Deal” video game based on a popular TV show, where contestants choose boxes with concealed cash amounts and negotiate with a banker to accept an offer or keep opening boxes. The aim is to secure the highest prize, avoiding smaller sums. Once a box is opened, its prize is shown and removed from potential winnings. Essentially, players choose between risky lotteries and certain cash offers. They also may exchange their box for any unopened one in the game. The decision to swap the box or to stick to the original one enables us to study whether the participants behave according to EUT or to Cumulative Prospect Theory (CPT). A previous study [25] that used TV show data from three countries demonstrated that: (i) CPT players are more likely to stick to their original choice of the box due to the embedded assumption of loss aversion and (ii) EUT players are indifferent between swapping and sticking.
@@
We recruited 78 study participants (members of the general public over 18 years of age) who made a total of 1,698 game decisions, including 486 swap or stick decisions. The fact that 63% of decisions were stick decisions and 37% were swap decisions demonstrated a higher consistency with EUT than with CPT. However, Cumulative Distribution Function (CDF) of swap or stick decisions revealed an intriguing result: while EUT predicts an approximately equal split between swapping or sticking decisions (the straight dash-dotted lines in Fig. 4), the distribution of swap or stick decisions (the dotted curve in Fig. 4) cannot be described by EUT.
@@
Since the magnetisation dynamics depends on the static field $H_{0}$ (Fig. 1b), in Fig. 4 we plot a magnetisation reversal that is recast as $(M_{z}/M_{s}+1)/2$ for the sake of comparison with the experimental data. The resulting curve satisfactorily reproduces the experimental decision-making behaviour demonstrated by the players of the game. The values of $H_{0}$ corresponding to the data points of this curve are, from left to right: 15, 6, 8,75, 9,6, 9,95, 10, 10, 10.75, 11, 11.52, 11.99, 12.1 and 12.5 kOe.
@@
Figure 4: Top $x$ and right $y$ axes: Cumulative probability distribution function for sticking decisions in the video game. The dash-dotted straight lines denote the EUT propensity to stick. The dashed curve corresponds to the experimental data. Bottom $x$ and left $y$ axes: The circular markers denote the simulated magnetisation reversal. The thin solid line is the guide to the eye only.
@@
Discussion—The functions of a human brain underpinning the phenomena like PR can be comprehended in detail through neuroscience methods [26]. Nonetheless, understanding human behaviour is pivotal not just for neuroscience but also for fields of finance, politics, religion and sociology. Thus, there is a demand for phenomenological models that can forecast decisions at both individual and societal levels quickly and efficiently.
@@
These models should not necessarily be based on neurobiological insights, mirroring the human cognition (drawing parallels with biological brain processes might even be unhelpful [27]). The primary criteria for evaluating a model of decision-making should be the ability of the model to accurately explain and predict imprecise and noisy decisions. The model we introduce in this paper follows this philosophy, aligning with the ongoing work in decision theory where models are not necessarily capturing human brain functions [28].
@@
STNO has been a foundational component in neuromorphic hardware systems that emulate the functions of the human brain [29, 30]. The operational principles of STNO obey the laws of quantum mechanics [31], even though our model’s Eq. (1) is classical. Quantum-mechanical models have demonstrated superior capacity in depicting decision-making compared with classical models [20]. Therefore, leveraging a realistic STNO structure in our model hints at the potential of STNOs to act as a neuromorphic system for simulating human decision-making processes. The so-designed system can be integrated with a reservoir computing algorithm [32], resulting in a machine learning technique that has an enhanced learning efficacy compared with classical systems [33] and also can simulate the human decision-making.
@@
Other quantum-mechanical systems can replicate decision-making [34]. However, the magnetisation reversal in STNO naturally extends important models of opinion shifts in social networks based on spin direction reversal [16]. Even though those models adopted the spin concept from physics, the processes causing the spin changes were largely synthetic. Thus, integrating a genuine model of magnetisation reversal into the opinion shift model [16] will enhance the efficacy of the latter, enabling it to analyse diverse decision-making scenarios.
@@
Conclusions—We demonstrated that the psychological phenomenon of PR can be modelled using the physical phenomenon of magnetisation reversal. Using a spin transfer torque as the mechanism of magnetisation reversal, we revealed the ability of the magnetisation reversal model to capture complex individual’s decisions made in psychological experiments. We have also shown that our model extends beyond PR and encompasses a wide range of decision-making regularities, reaffirming the observation that human preferences are imprecise.
@@
Numerical method—We model the dynamics of magnetisation using Landau-Lifshitz-Gilbert equation [35]:
where $\gamma$ is the gyromagnetic ratio. The first term of the right-hand side of Eq. (1) governs the precession of $\bf M$ (Fig. 1a) about the direction of the effective magnetic field $\textbf{H}_{eff}=H_{0}\textbf{e}_{\rm{z}}+\textbf{H}+\textbf{H}_{ex}$, where $H_{0}$ is the external static magnetic field orientated along the z-axis, H is the dynamic field due to currents and magnetic sources such as demagnetising field and eddy current fields and $\textbf{H}_{ex}$ is the exchange field [36]. The dissipative torque is [35]
where $M_{s}$ is the saturation magnetisation of the “free” magnetisation layer and $\alpha_{G}$ is Gilbert damping parameter [36, 35]. The Slonczewski-Berger torque is
where $I$ is the strength and $\hat{e}_{p}$ is the direction of the spin polarisation of the current. The parameter $\sigma_{0}$ incorporates fundamental physical constants and the thickness of the “free” magnetisation layer [35]. Equation (1) is solved consistently with Maxwell’s equations using a finite-difference time-domain (FDTD) method and material parameters presented in [36]. We model a point-contact STNO structure experimentally studied in [24]. Since we employ a one-dimensional FDTD scheme, the strength of the current $I$ required to produce a magnetisation reversal in our model is lower than that in [24].

### v2

@@ 2.3 Results
In Fig. 1b we plot the dynamics of the $M_{z}$ component of the magnetisation vector ${\bf M}$ in the “free” layer as a function of the electric current strength for different value of the applied magnetic field $H_{0}$. The trajectories of ${\bf M}$ for $H_{0}=8$ kOe computed at the simulated time of 1 ns are shown in Fig. 5b–c. Figure 1b demonstrates that the magnetisation dynamics correctly approximates the sigmoid probabilistic PR curve in Fig. 1a but the strength of the driving electric current plays the role of parameter $A$.
The experimental curves in Fig. 4 can be regarded as the particular cases of the sigmoid function and they can be simulated using the magnetisation reversal observed at different values of $H_{0}$ (for example, compare the experimental CDF curve in Fig. 4c.ii with the magnetisation reversal curve for 9 kOe in Fig. 1b). We also note that it is computationally straightforward to use our model to approximate curves whose shape deviates from a sigmoid function. To demonstrate this possibility, in Study 2 we unleash the potential of the magnetisation reversal model applying it to the analysis of a more complex experimental dataset.
3 Study 2
@@
We further validate our model using a dataset obtained from a “Deal or No Deal” video game based on a popular TV show, where contestants choose boxes with concealed cash amounts and negotiate with a banker to accept an offer or keep opening boxes. The aim is to secure the highest prize, avoiding smaller sums. Once a box is opened, its prize is shown and removed from potential winnings. Essentially, players choose between risky lotteries and certain cash offers. They also may exchange their box for any unopened one in the game. The decision to swap the box or to stick to the original one enables us to study whether the participants behave according to EUT or to Cumulative Prospect Theory (CPT). A previous study blavatskyy2010endowment that used TV show data from three countries demonstrated that: (i) CPT players are more likely to stick to their original choice of the box due to the embedded assumption of loss aversion and (ii) EUT players are indifferent between swapping and sticking.
@@
We recruited 78 study participants (members of the general public over 18 years of age, informed consent was obtained from all participants as per the procedure used in Study 1) who made a total of 1,698 game decisions, including 486 swap or stick decisions. The fact that 63% of decisions were stick decisions and 37% were swap decisions demonstrated a higher consistency with EUT than with CPT. However, Cumulative Distribution Function (CDF) of swap or stick decisions revealed an intriguing result: while EUT (the straight dashed lines in Fig. 6) predicts an approximately equal split between swapping or sticking decisions, the distribution of swap or stick decisions (the dotted curve in Fig. 6) cannot be described by EUT.
@@
Since the magnetisation dynamics depends on the static field $H_{0}$ (Fig. 1b), in Fig. 6 we plot a magnetisation reversal that is recast as $(M_{z}/M_{s}+1)/2$ for the sake of comparison with the experimental data. The resulting curve satisfactorily reproduces the experimental decision-making behaviour demonstrated by the players of the game. The values of $H_{0}$ corresponding to the data points of this curve are, from left to right: 15, 6, 8.75, 9.6, 9.95, 10, 10, 10.75, 11, 11.52, 11.99, 12.1 and 12.5 kOe.
@@
Figure 6: Top $x$ and right $y$ axes: Cumulative probability distribution function for sticking decisions in the video game. The dashed straight lines denote the EUT propensity to stick. The dotted curve corresponds to the experimental data. Bottom $x$ and left $y$ axes: The circular markers denote the simulated magnetisation reversal. The thin solid line is the guide to the eye only.
4 Discussion
@@
The functions of a human brain underpinning the phenomena like PR can be comprehended in detail through neuroscience methods Sha21 . Nonetheless, understanding human behaviour is pivotal not just for neuroscience but also for fields of finance, politics, religion and sociology. Thus, there is a demand for phenomenological models that can forecast decisions at both individual and societal levels quickly and efficiently.
@@
These models should not necessarily be based on neurobiological insights, mirroring the human cognition (drawing parallels with biological brain processes might even be unhelpful Tad16 ). The primary criteria for evaluating a model of decision-making should be the ability of the model to accurately explain and predict imprecise and noisy decisions. The model we introduce in this paper follows this philosophy, aligning with the ongoing work in decision theory where models are not necessarily capturing human brain functions wakker2010prospect .
Importantly, our model is not driven by experimental datasets. In fact, the psychological data presented in this paper are needed only to validate the model. The validation process requires relatively compact datasets and its main aim is to establish a range of model parameters that can be used in further studies. Once suitable parameters have been established, the model can produce new results independently of experimental data.
This approach has certain limitations, including the need to adjust the model parameters to analyse a previously unconsidered social or economical scenario (a more comprehensive discussion of the perceived limitations of the proposed model is given in Sect. 5). However, the adjustment process should be relatively simple since it requires one to validate the model against another compact set of experimental data. Overall, results produced by a properly adjusted model can help reduce the need to conduct complex experiments involving many human participants.
@@
STNO has been a foundational component in neuromorphic computing hardware systems that emulate the functions of the human brain Fur18 ; Rio19 . The operational principles of STNO obey the laws of quantum mechanics Bra22 , even though our model’s Eq. (1) is classical. Quantum-mechanical models have demonstrated superior capacity in depicting decision-making compared with classical models Bus12 . Therefore, leveraging a realistic STNO structure in our model hints at the potential of STNOs to act as a neuromorphic system for simulating human decision-making processes. The so-designed system can be integrated with a reservoir computing algorithm Mak23_review , resulting in a machine learning technique that has an enhanced learning efficacy compared with classical systems Dud23 and also can simulate the human decision-making.
@@
Other quantum-mechanical systems can replicate decision-making Nar14 . However, the magnetisation reversal in STNO naturally extends important models of opinion shifts in social networks based on spin direction reversal Red19 . Even though those models adopted the spin concept from physics, the processes causing the spin changes were largely synthetic. Thus, integrating a genuine model of magnetisation reversal into the opinion shift model Red19 will enhance the efficacy of the latter, enabling it to analyse diverse decision-making scenarios.
5 Future Research
The current work is a plausible attempt to phenomenologically describe psychological and decision-making phenomena using a purely physical model. Although this approach is mathematically similar to models belonging to the quantum decision theory group, our model was developed independently, exploiting the ability of classical physics to successfully explain certain quantum-mechanical phenomena.
Admittedly, our model cannot capture the complexity of cognitive processes in full. However, creating such a model goes well beyond the scope of this paper. Yet, as famously written by Box and Draper, the approximate nature of the model must always be borne in mind Box87 . In other words, any model has its advantages and disadvantages and its true performance can be demonstrated only in tests involving real-life data. Supporting this point of view and recognising the benefits of a comparison of different models Bus00 , in this paper we establish the ground truth by benchmarking our model against large datasets obtained in diverse experimental scenarios, which a standard research approach in the fields of physics and engineering. It is also likely that additional limitations of our model will be revealed at later stages, when the model will be used by different users. This process is natural and it will help us and other researchers following our work improve the model.
Importantly, since we aimed to make our model accessible to a broad range of non-experts in physics, we deliberately kept the model equations as simple as possible. Hence, future research could explore more intricate physical processes and more advanced quantum-physical models that may offer a closer analogue to the multifaceted nature of human cognition and decision-making. Indeed, the experimental setups, such as those employed in our “Deal or No Deal” study, provide valuable insights but also come with limitations regarding, for example, the range of decision-making scenarios they can simulate. Further studies could employ more diverse and real-life decision-making contexts, enhancing the applicability and robustness of our findings.
This study predominantly focuses on an external, observable behaviour perspective. Integrating other kind of insights, especially from fields like macroeconomics, could provide a more holistic understanding of decision-making processes, bridging the gap between physical models and the neurological underpinnings of human behaviour. The inclusion of these factors in our model should not represent significant technical difficulties. Indeed, strictly speaking, if a decision-making process can be represented by a continuous function at least in principle then our model or its modifications should be able to approximate it with a reasonable accuracy. Yet, our physical model can naturally incorporate noise, interference and dissipation, all defined physically, which are the processes that were previously explored in models of cognition and decision-making Khr06 ; Aer22 ; Aer22_1 .
The computational models employed in psychology, while effective, may not fully capture the dynamism and adaptability of human decision-making. Future research could incorporate machine learning and AI to develop more sophisticated models that can dynamically adapt and learn from new data, mirroring the evolving nature of human preferences. One example of a relevant study is our recent work on a quantum-inspired neural network model of optical illusions that exploits a chaotic physical system as a network Mak24 . The cited work aligns with a broader collective effort to create neural networks that could predict complex socio-economic processes using the laws of physics (for a review of methodological aspects see, e.g., Mak23_review ).
Our study primarily considers decision-making from a psychological and physical perspective, potentially overlooking the influence of socio-cultural factors, which may be of interest. Subsequent research could integrate these variables, examining how cultural, social, and environmental factors interplay with the cognitive processes in decision-making. We are currently working in this direction and the results of that project will be reported elsewhere.
By addressing these limitations, future research can not only refine and expand upon our current findings but also contribute to a more nuanced and comprehensive understanding of human decision-making processes. More specifically, the current research can lead to a deeper integration of physical sciences with behavioural economics and psychology. Future research could contribute by developing comprehensive models that further incorporate principles from quantum mechanics, statistical physics and complexity science to capture the nuanced aspects of human decision-making, as exemplified by neural network models empowered by quantum technologies Mak24 . There is also an opportunity to explore the neurobiological underpinnings of the decision-making phenomena modelled by physical systems. Indeed, future studies could use neuroimaging techniques to validate the predictions of physical models against observed brain activity during decision-making tasks (see, e.g., the works on artificial neural network that mimic the operation of a biological brain Rao22 and relevant works on physics-informed machine learning models Nak22 ; Mak23_review ).
Finally, advanced computational techniques like machine learning and AI could be employed to refine the models based on large datasets. In this context, future research could focus on the development of algorithms that learn from behavioural data to predict decision-making patterns and preference changes over time. A number of papers developing relevant ideas have already been published Bus17 ; Khr18 ; Khr20 ; Sal23 . Future research could utilise such models to simulate the effects of policy changes on public behaviour: by understanding how individuals might react to different policy interventions more effective and targeted strategies can be developed. Indeed, there have been several demonstrations of the possible application of quantum-mechanics in political studies, especially in the realm of international relations studies Jak15 ; All18 . Importantly, the cited works present rather a general philosophical discussion of such models, thereby highlighting unexplored fields of research where physical model of decision-making can find specific practical applications.
6 Conclusions
@@
We demonstrated that the psychological phenomenon of PR can be modelled using the physical phenomenon of magnetisation reversal. Using a spin transfer torque as the mechanism of magnetisation reversal, we revealed the ability of the magnetisation reversal model to capture complex individual’s decisions made in psychological experiments. We have also shown that our model extends beyond PR and encompasses a wide range of decision-making regularities, reaffirming the observation that human preferences are imprecise.
6.1 Implications on Literature
This research contributes to a dynamic and ever-evolving body of literature that challenges traditional economic and psychological theories. It encourages the development of new theories that accommodate the stochastic and probabilistic nature of human decision-making. The use of physical models to explain psychological phenomena can inspire new research methodologies that are more robust and predictive of real-world behaviours. The findings demonstrate the value of cross-disciplinary research and can stimulate collaborative studies that bridge gaps between fields such as cognitive psychology, economics, physics and computer science.
6.2 Implications on Practice and Policy
In practice, these models can inform the design of behavioural interventions that are more aligned with how people actually make decisions, potentially increasing their efficacy. The ability to more accurately predict human behaviour can revolutionise economic forecasting and financial planning, providing tools for better risk assessment and market analysis. Policymakers can use insights from this research to craft policies that consider the probabilistic nature of human decisions, leading to more nuanced and effective public policies. The interdisciplinary nature of this research can also influence educational curricula, promoting a more integrative approach to teaching economics, psychology and physics.
@@ 6.2 Implications on Practice and Policy
6.3 Implications on Interdisciplinary Research in Physics
Last but not least, the present work contributes to a more than four decades long endeavour to establish the field of sociophysics Gal22 . Although sociophysics has maturated as an independent discipline, it has been unable to find a secure place in specialised physical journals and university classrooms yet. The present work demonstrates that a physical model developed to investigate the fundamental physical effect of magnetisation reversal can be applied in the domain of social sciences without the need to modify the model parameters. While some physicists may find this result to be contradictory, our findings demonstrate that the actual frontiers of physics are located well beyond our current understanding and that the point where a physical topic becomes a non-physical one is more difficult to define than is immediately apparent.
Ethics statement
This work was approved by Humanities and Social Sciences Research Ethics Committee (HSSREC) of the University of Warwick (approval number 31/13-14). Upon arrival to the laboratory, each participant was provided with the study information sheet and a consent form. Informed consent was obtained from all participants in the study. All participants were adults (18 years of age or older). No minors participated in the experiments; therefore, obtaining parental/guardian consent was not applicable to this work.
Data Availability
All data relevant to this work are presented in this paper. Large-size relevant raw data are uploaded to GitHub: https://github.com/BehaviouralDataScience/Magnetisation-Reversal-Model alongside the software used to process them.
CRediT authorship contribution statement
Ivan S. Maksymov: Writing–original draft, Writing–review & editing, Visualization, Methodology, Investigation, Formal analysis, Conceptualization.
Ganna Pogrebna: Writing–original draft, Writing–review & editing, Visualization, Project administration, Methodology, Investigation, Formal analysis, Conceptualization.
All authors gave final approval for publication and agreed to be held accountable for the work performed therein.
Competing interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of AI use
We have not used AI-assisted technologies in creating this article.
Funding
We received no funding for this study.
Acknowledgements
Not applicable.